OpenStudy (anonymous):

Find the intercepts and then graph the line 3x-4y=18

OpenStudy (anonymous):

x-intercept: (6,0) y-intercept: (0,−9/2)

OpenStudy (anonymous):

need explanation?

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

Multiply −4 by each term inside the parentheses. More Detail Multiply −4 by the 0 inside the parentheses. Multiply −4 by 0 to get 0. Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. Divide each term in the equation by 3. Simplify the left-hand side of the equation by canceling the common factors. Cancel the common factor of 3 in 3x/3. Reduce the expression 3x/3 by removing a factor of 3 from the numerator and denominator. x=18/3 Simplify the equation. Cancel the common factor of 3 in 18/3. Reduce the expression 18/3 by removing a factor of 3 from the numerator and denominator. x=6 To find the y-intercept, substitute in 0 for x and solve for y. Multiply 3 by each term inside the parentheses. Multiply 3 by the 0 inside the parentheses. Multiply 3 by 0 to get 0. Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. Divide each term in the equation by −4. Simplify the left-hand side of the equation by canceling the common factors. Move the minus sign from the denominator to the front of the expression. Cancel the common factor of 4 the expression 4y/4. Reduce the expression 4y4 by removing a factor of 4 from the numerator and denominator. y=18/−4 Solve the equation. Move the minus sign from the denominator to the front of the expression. y=−18/4 Cancel the common factor of 2 the expression −18/4. Reduce the expression −18/4 by removing a factor of 2 from the numerator and denominator. y=−9/2 These are the x and y intercepts of the equation 3x−4y=18. x-intercept: (6,0) y-intercept: (0,−92)

OpenStudy (anonymous):

-9/2 y itnercept

OpenStudy (anonymous):

but why did graph line go throug -4 on y intercept

OpenStudy (anonymous):

but the answer for y is -9/2

OpenStudy (anonymous):

cause -9/2 also equals = -4 and 1/2

OpenStudy (anonymous):

orless

OpenStudy (anonymous):

how many times does 2 go into 9? 4 times with 1/2 left right

OpenStudy (anonymous):

oh good thanks goog nite

OpenStudy (anonymous):

good night